An examination paper which is divided into 2 groups consisting of 3 and 4 questions respectively, carries the note:"It is not necessary to answer all questions. One question must be answered from each group."In how many ways can an examinee select the questions?

Grade:

1 Answers

abhay tiwari
12 years ago

number of way can group1 question be selected =3C1

number of way cab group2 question be selected =4C1

numeber of way can both be selected simultaneously =3C1*4C1=4*3=12 ANSWER